OpenStudy (anonymous):

evluate ∫ (x+1) / x^2 (x−1) ԁx

OpenStudy (dumbcow):

partial fractions \[\frac{x+1}{x^2 (x-1)} = \frac {Ax +B}{x^2} + \frac{C}{x-1}\] \[(Ax+B)(x-1) + Cx^2 = x+1\] \[A +C = 0\] \[B -A =1\] \[-B = 1\] solve for A,B,C